Determine if the following statements are true or false. Show work please.


  1. 4/15=2/7                                   2. 4/6=10/15
1
Expert's answer
2021-08-06T08:26:08-0400
  1. 4/15=2/7    FALSE because the two fractions cannot be simplified further, and they are not equal.
  2. 4/6=10/15 TRUE because 4/6 = 2/3, 10/15 if we simplify using 5 we get 2/3. 2/3=2/3 hence it is TRUE